A photon and an electron each have an energy of .eV100.63 What are their wavelengths? Which of these would you use to probe atomic structures?
1
Expert's answer
2020-03-25T11:00:03-0400

Suppose that "E=100.63 keV"


For electron


"\lambda_e=\frac{h}{mv}=\frac{h}{\sqrt{{2mE}}}=\frac{6.62\cdot10^{-34}}{\sqrt{{2\cdot 9.11\cdot10^{-31}\cdot 1.6\cdot10^{-19}\cdot100630}}}="


"=3.86\cdot10^{-12}m=3.86pm"


For proton


"\lambda_p=\frac{h}{mv}=\frac{h}{\sqrt{{2mE}}}=\frac{6.62\cdot10^{-34}}{\sqrt{{2\cdot 1.67\cdot10^{-27}\cdot 1.6\cdot10^{-19}\cdot100630}}}="


"=9.03\cdot10^{-14}m\approx0.09pm"


I would use electrons to probe atomic structures.